# ]https://zf-wei.github.io

### v1

@@ III.2 Random walk methods
Random walk methods learns network embeddings of graph nodes by modeling a stream of short random walks. We chose DeepWalk [31], LINE [32], and node2vec [33], as widely used methods of this family.
@@ III.2.1 DeepWalk
DeepWalk [31] extends language modeling techniques to graphs, departing from words and sentences. This algorithm leverages local information acquired through random walks, treating these walks as analogous to sentences in the word2vec [65] language modeling approach. To generate a random walk originating from a specified starting node, neighbors of the current node in the walk are randomly selected and added to the walk iteratively until the intended walk length is achieved. In our experiments, we use the following parameters: random walk length $t=40$, window size $w=10$, number of walks per node $n=80$.
Our experimental results on the impact of edge removal, generated through the DeepWalk embedding method, are reported in Figs. 14 and 15. We can observe that networks with stronger initial partitions (i.e., lower $\mu$ values) experience less perturbation impact and targeted node selection has a significant influence on clusters. Examining heatmaps in Fig. 16(a), it is evident that for strongly clustered networks, the DeepWalk embedding method produces higher ECS scores for smaller network sizes and lower embedding dimensions.
Figure 14: Mean ECS using DeepWalk method over the percentage of selected nodes on LFR benchmark graphs with $1,000$ nodes. (a) and (b) use random node selection while (c) and (d) use targeted node selection.
Figure 15: Mean ECS using DeepWalk method over the percentage of selected nodes on LFR benchmark graphs with $10,000$ nodes. (a) and (b) use random node selection while (c) and (d) use targeted node selection.
@@ III.2.2 Large-scale information network embedding (LINE)
The large-scale information network embedding (LINE) method aims to position nodes in close proximity to each other as their similarity increases. LINE can be implemented based on first-order similarity, second-order similarity, or both. According to [14], it works the best when implemented based on first-order similarity, i.e., adjacency of nodes. In fact, for un-weighted network, LINE method based on first-order similarity minimize this objective function:
@@
The optimization process of LINE method utilizes stochastic gradient descent, which is enhanced by an edge-sampling treatment, as detailed in [32]. Specifically, the utilization of the LINE embedding method necessitates the presence of GPU for computation. In our study, the machine learning model was trained using a batch size of $8192$ over $100$ epochs, facilitating iterative updates to the embedding vector across the entire training dataset.
Our experimental results, generated through the LINE embedding method, are presented in Figs. 17 and 18. The LINE embedding method produces higher ECS scores when applied to larger networks. Our analysis of networks with 1,000 nodes in Fig. 17 reveals that curves exhibit a steep decline followed by a subsequent increase, but we notice that increasing the embedding dimensions in LINE is expected to mitigate the sharpness of this decline and increase; see Appendix A for details. Owing to the substantial computational demands of the LINE embedding method, we omit the heatmap analysis specifically for LINE.
@@
Figure 17: Mean ECS using LINE method over the percentage of selected nodes on LFR benchmark graphs with $1,000$ nodes. (a) and (b) use random node selection while (c) and (d) use targeted node selection.
Figure 18: Mean ECS using LINE method over the percentage of selected nodes on LFR benchmark graphs with $10,000$ nodes. (a) and (b) use random node selection while (c) and (d) use targeted node selection.
@@
node2vec employs a similar optimization procedure to DeepWalk, but the process of generating “sentences” differs [33]. Specifically, simple random walk is used for DeepWalk, while biased random walk is utilized for node2vec. These biased random walks are composed of a blend of steps following both breadth-first and depth-first search strategies, with parameters $p$ and $q$ controlling the respective influences of these two strategies (see [33] for details about how random walk are generated in node2vec). Therefore, node2vec produces higher-quality and more informative embeddings than DeepWalk. We use the following parameters in our experiments: random walk length $t=40$, window size $w=10$, number of walks per node $n=80$, biased walk weights $p=1$ and $q=1$. Our experimental results on the impact of edge removal, generated through the node2vector embedding method, are reported in Figs. 19 and 20. The heatmaps in Fig. 21 illustrate that node2vec proves to be robust method to produce community detection outcome with high ECS, particularly for strongly clustered networks. In Fig. 21(b) for weakly clustered networks (i.e., $\mu=0.5$), to get high ECS, in general, lower embedding dimensions produce more robust community partition for networks with small sizes; higher embedding dimensions produce more robust community partition for networks with larger sizes.
Figure 19: Mean ECS using node2vec method over the percentage of selected nodes on LFR benchmark graphs with $1,000$ nodes. (a) and (b) use random node selection while (c) and (d) use targeted node selection.
Figure 20: Mean ECS using node2vec method over the percentage of selected nodes on LFR benchmark graphs with $10,000$ nodes. (a) and (b) use random node selection while (c) and (d) use targeted node selection.
@@
When considering the application of random walk embeddings for community detection, node2vec consistently proves to be a more robust choice over DeepWalk and LINE, across networks of varying sizes, the strength of the initial community partition, and perturbation type.
@@ III.4 Results on real-world networks
In this section, we present the results of our experiments conducted on two real-world networks with labeled communities, i.e., the email-EU-core network and the AS network. The email-EU-core network is constructed using email data sourced from a prominent European research institution, comprising $986$ nodes and $16,687$ edges [63]. The AS network, derived from the AS Internet topology data collected in June 2009 by the Archipelago active measurement infrastructure, represents the inter-domain Internet topology with Autonomous Systems (ASes) as nodes and AS peerings as links, forming an AS-level topology graph. This dataset comprises $23,748$ nodes and $58,414$ edges. We select them because they closely resemble power-law degree and community size distributions, as the synthetic networks counterpart (i.e., LFR benchmark graphs). To confirm the alignment of these real-world networks with the properties of LFR networks, we conducted statistical tests using the Python package powerlaw based on the statistical test derived in the work by Clauset et al [66]. The results indicate that both the email-EU-core network and the AS network exhibit power-law characteristics [67]. Specifically, in the email-EU-core network, the degree distribution adheres to a power-law distribution and the community size sequence follows exponential distribution; while in the AS network, both degree and community size distributions conform to power laws.
@@
We report experimental results using different embedding methods on these two real-world networks in Figs. 23 and 24. Comparing the two rows of Fig. 23, we can see that when employing targeted node selection, the ECS experiences a more significant decrease than the case of random node selection. For example, for the email-EU-core network, when perturbing it at $50\%$ and using node2vec with $16$-dimensional embedding, we notice that targeted node selection decreases ECS from $0.41$ to $0.31$ (a decrease of $24.3\%$, see Fig. 23(c)), but ECS is only decreased by around $0.02$ by random node selection. Furthermore, LLE and node2vec proved to be the superior choices among all seven embedding methods. In fact, curves corresponding to these two methods in Figs. 23 and 24 are higher in position than other curves.
Figure 23: Mean ECS for various methods over the percentage of selected nodes on email-EU-core network (with $986$ nodes). (a) and (b) use random node selection while (c) and (d) use targeted node selection. (a) and (c) corresponds to $16$-dimensional embeddings while (b) and (d) corresponds to $32$-dimensional embeddings.
Figure 24: Mean ECS for various methods over the percentage of selected nodes on AS network (with $23,748$ nodes). (a) and (b) use random node selection while (c) and (d) use targeted node selection. (a) and (c) corresponds to $16$-dimensional embeddings while (b) and (d) corresponds to $32$-dimensional embeddings.
@@ IV Conclusion
Appendix A LINE method with higher embedding dimensions
For the LINE embedding method, our analysis of networks with 1,000 nodes in Fig. 17 reveals a distinct trend: the curves exhibit a steep decline followed by a subsequent increase. According to our experiments, increasing the embedding dimensions in LINE is expected to mitigate the sharpness of this decline and increase, rendering the pattern less pronounced, as indicated in Fig. 25.
Figure 25: Mean ECS using LINE method over the percentage of selected nodes uniformly at random on LFR benchmark graphs with $1,000$ nodes. The first plot displays results using $128$-dimensional embedding; the second plot displays results using $256$-dimensional embedding.
However, for other embedding methods, a higher embedding dimension doesn’t necessarily lead to better performance with higher ECS scores. For instance, M-NMF with higher dimensions results in lower ECS scores, as shown in Fig. 26, compared to the first row of Fig. 11.
@@ Appendix A LINE method with higher embedding dimensions
Figure 26: Mean ECS using M-NMF method over the percentage of selected nodes uniformly at random on LFR benchmark graphs with $1,000$ nodes. The first plot displays results using $128$-dimensional embedding; the second plot displays results using $256$-dimensional embedding.
@@ Acknowledgements.
This paper has been authored by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. The publisher, by accepting the article for publication, acknowledges that the U.S. government retains a nonexclusive, paid up, irrevocable, world-wide license to publish or reproduce the published form of the manuscript, or allow others to do so, for U.S. government purposes. The DOE will provide public access to these results in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan). This research was supported in part by an appointment with the National Science Foundation (NSF) Mathematical Sciences Graduate Internship (MSGI) Program. This program is administered by the Oak Ridge Institute for Science and Education (ORISE) through an interagency agreement between the U.S. Department of Energy (DOE) and NSF. ORISE is managed for DOE by ORAU. All opinions expressed in this paper are the author’s and do not necessarily reflect the policies and views of NSF, ORAU/ORISE, or DOE. This research was supported in part by an appointment to the Oak Ridge National Laboratory GRO Program, sponsored by the U.S. Department of Energy and administered by the Oak Ridge Institute for Science and Education. This research was supported in part by Lilly Endowment, Inc., through its support for the Indiana University Pervasive Technology Institute.

### v2

@@ III.2 Random walk methods
Random walk methods learn network embeddings of graph nodes by modeling a stream of short random walks. We chose DeepWalk [31], LINE [32], and node2vec [33], as widely used methods of this family.
@@ III.2.1 DeepWalk
DeepWalk [31] extends language modeling techniques to graphs, departing from words and sentences. This algorithm leverages local information acquired through random walks, treating these walks as analogous to sentences in the word2vec [65] language modeling approach. To generate a random walk originating from a specified starting node, neighbors of the current node in the walk are randomly selected and added to the walk iteratively until the intended walk length is achieved.
In our experiments, we use the following parameters: random walk length $t=40$, window size $w=10$, number of walks per node $n=80$.
Our experimental results on the impact of edge removal, generated through the DeepWalk embedding method with dimension $16$, are reported in Fig. 10. Our experimental results using DeepWalk embedding method with dimension $32$ are collected in Fig. 21. We can observe that networks with stronger initial partitions (i.e., lower $\mu$ values) experience less perturbation impact and targeted node selection has a significant influence on clusters. Examining heatmaps in Fig. 11(a), it is evident that for strongly clustered networks, the DeepWalk embedding method produces higher ECS scores for smaller network sizes and lower embedding dimensions.
Figure 10: Mean ECS using DeepWalk method over the percentage of selected nodes on LFR benchmark graphs with $16$-dimensional embeddings.Random selection is employed in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) represent networks with $1,000$ nodes, and (b) and (d) represent those with $10,000$ nodes.
@@ III.2.2 Large-scale information network embedding (LINE)
The large-scale information network embedding (LINE) method aims to position nodes in close proximity to each other as their similarity increases. LINE can be implemented based on first-order similarity, second-order similarity, or both. According to [14], it works the best when implemented based on first-order similarity, i.e., adjacency of nodes. In fact, for unweighted networks, the LINE method based on first-order similarity minimizes this objective function:
@@
The optimization process of the LINE method utilizes stochastic gradient descent, which is enhanced by an edge-sampling treatment, as detailed in [32]. Specifically, the utilization of the LINE embedding method necessitates the presence of GPU for computation. In our study, the machine learning model was trained using a batch size of $8192$ over $100$ epochs, facilitating iterative updates to the embedding vector across the entire training dataset.
Our experimental results generated through the LINE embedding method with dimension $16$ are presented in Fig. 12. Our experimental results using LINE embedding method with dimension $32$ are collected in Fig. 22 (see Appendix A). The LINE embedding method produces higher ECS scores when applied to larger networks. Our analysis of networks with 1,000 nodes in the first column of Fig. 12 reveals that curves exhibit a steep decline followed by a subsequent increase, but we notice that increasing the embedding dimensions in LINE is expected to mitigate the sharpness of this decline and increase; see Appendix B for details. Owing to the substantial computational demands of the LINE embedding method, we omit the heatmap analysis specifically for LINE.
@@
Figure 12: Mean ECS using LINE method over the percentage of selected nodes on LFR benchmark graphs with $16$-dimensional embeddings. (a) and (b) follow random node selection, whereas (c) and (d) adopt targeted node selection. Networks with $1,000$ nodes correspond to (a) and (c), and those with $10,000$ nodes correspond to (b) and (d).
@@
node2vec employs a similar optimization procedure to DeepWalk, but the process of generating “sentences” differs [33]. Specifically, simple random walk is used for DeepWalk, while biased random walk is utilized for node2vec. These biased random walks are composed of a blend of steps following both breadth-first and depth-first search strategies, with parameters $p$ and $q$ controlling the respective influences of these two strategies (see [33] for details about how random walk are generated in node2vec). Therefore, node2vec produces higher-quality and more informative embeddings than DeepWalk.
We use the following parameters in our experiments: random walk length $t=40$, window size $w=10$, number of walks per node $n=80$, biased walk weights $p=1$ and $q=1$.
Our experimental results on the impact of edge removal, generated through the node2vec embedding method with dimension $16$, are reported in Fig. 13. Our experimental results using node2vec embedding method with dimension $32$ are collected in Fig. 23 (see Appendix A). The heatmaps in Fig. 14 illustrate that node2vec proves to be a robust method to produce community detection outcomes with high ECS, particularly for strongly clustered networks. In Fig. 14(b) for weakly clustered networks (i.e., $\mu=0.5$), to get high ECS, in general, lower embedding dimensions produce more robust community partition for networks with small sizes; higher embedding dimensions produce more robust community partition for networks with larger sizes.
Figure 13: Mean ECS using node2vec method over the percentage of selected nodes on LFR benchmark graphs with $16$-dimensional embeddings. Random node selection is applied in (a) and (b), while (c) and (d) use targeted selection. (a) and (c) deal with $1,000$-node networks, and (b) and (d) with $10,000$-node networks.
@@ III.2.4 Summary of random walk methods
From our discussion of each method, we observe that DeepWalk may exhibit an unexpected increase while LINE experiences unexpected oscillations. In contrast, node2vec is free from these issues and consistently achieves the highest ECS similarity scores. When considering the application of random walk embeddings for community detection, node2vec consistently proves to be a more robust choice over DeepWalk and LINE, across networks of varying sizes, the strength of the initial community partition, and perturbation type.
@@ III.4 Results on real-world networks
In this section, we present the results of our experiments conducted on two real-world networks with labeled communities, i.e., the email-EU-core network and the AS network. The email-EU-core network is constructed using email data sourced from a prominent European research institution, comprising $986$ nodes and $16,687$ edges [63]. The AS network, derived from the AS Internet topology data collected in June 2009 by the Archipelago active measurement infrastructure, represents the inter-domain Internet topology with Autonomous Systems (ASes) as nodes and AS peerings as links, forming an AS-level topology graph. This dataset comprises $23,748$ nodes and $58,414$ edges. We select them because they closely resemble power-law degree and community size distributions, as the synthetic networks counterpart (i.e., LFR benchmark graphs). To confirm the alignment of these real-world networks with the properties of LFR networks, we conducted statistical tests using the Python package powerlaw based on the statistical test derived in the work by Clauset et al [66]. The results indicate that both the email-EU-core network and the AS network exhibit power-law characteristics [67]. Specifically, in the email-EU-core network, the degree distribution adheres to a power-law distribution and the community size sequence follows an exponential distribution; while in the AS network, both degree and community size distributions conform to power laws.
@@
We report experimental results using different methods with $16$-dimensional embedding on these two real-world networks in Fig. 16. Comparing the two rows of Fig. 16, we can see that when employing targeted node selection, the ECS experiences a more significant decrease than the case of random node selection. For example, for the email-EU-core network, when perturbing it at $50\%$ and using node2vec with $16$-dimensional embedding, we notice that targeted node selection decreases ECS from $0.41$ to $0.31$ (a decrease of $24.3\%$, see Fig. 16(c)), but ECS is only decreased by around $0.02$ by random node selection (Fig. 16(a)). Furthermore, LLE and node2vec proved to be the superior choices among all seven embedding methods. In fact, curves corresponding to these two methods in Fig. 16 are higher in position than other curves in general. The experimental results using different methods with $32$-dimensional embedding on the two real-world networks are depicted in Fig. 24.
Figure 16: Mean ECS for various methods over the percentage of selected nodes on the two real-world networks with $16$-dimensional embeddings. (a) and (b) use random node selection while (c) and (d) use targeted node selection. (a) and (c) corresponds to the email-EU-core network while (b) and (d) corresponds to the AS network.
@@ IV Conclusion
Appendix A Experimental results using $32$-dimensional embedding
In this appendix, we present figures on experimental results using embedding dimension $32$. In Figs. 17, 18, 19, 20, 21, 22 and 23, subfigures in the top row, i.e., (a) and (b), correspond to random node selection. More specifically, subfigure (a) corresponds to experimental results on networks with $1,000$ nodes and subfigure (b) depicts experimental results on networks with $10,000$ nodes. Similarly, subfigures in the bottom row, i.e., (c) and (d), correspond to targeted node selection. More specifically, subfigure (c) corresponds to experimental results on networks with $1,000$ nodes and subfigure (d) depicts experimental results on networks with $10,000$ nodes.
We notice that some embedding methods exhibit different robustness with varying embedding dimensions. For instance, when using the Laplacian eigenmap method, in networks with strong community partitions (i.e., $\mu=0.01$ or $\mu=0.1$), the embedding dimension significantly impacts the decay rate of similarity, regardless of the node selection strategy. In Fig. 2(a), with an embedding dimension of $16$ and random node selection, the curve for $\mu=0.1$ remains stable around an ECS value of approximately $1.0$ until $60\%$ of nodes are selected. Between $60\%$ and $85\%$ of nodes selected, the curve gradually decreases from $1.0$ to $0.84$ (a $16\%$ decrease). In contrast, in Fig. 17(a), with a $32$-dimensional embedding, the curve for $\mu=0.1$ remains stable around $1.0$ only until $25\%$ of nodes are selected. Then, between $25\%$ and $85\%$ of nodes selected, the curve sharply decreases from $1.0$ to $0.53$ (a $47\%$ decrease). For the HOPE embedding method, our results suggest that a $32$-dimensional embedding (i.e., Fig. 19) yields higher ECS scores than a $16$-dimensional embedding (i.e., Fig. 6). Specifically, the data points on the green curve (corresponding to $\mu=0.4$) in Fig. 19(a) have higher ECS scores than those on the green curve in Fig. 6(a) at respective percentages of selected nodes.
@@ Appendix A Experimental results using $32$-dimensional embedding
Figure 17: Mean ECS using LE method over the percentage of selected nodes on LFR benchmark graphs with $32$-dimensional embeddings. Random node selection is used in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) are for networks with $1,000$ nodes, and (b) and (d) are for networks with $10,000$ nodes.
Figure 18: Mean ECS using LLE method over the percentage of selected nodes on LFR benchmark graphs with $32$-dimensional embeddings. (a) and (b) apply random node selection, whereas (c) and (d) use targeted selection. Networks in (a) and (c) have $1,000$ nodes, and those in (b) and (d) have $10,000$ nodes.
Figure 19: Mean ECS using HOPE method over the percentage of selected nodes on LFR benchmark graphs with $32$-dimensional embeddings. While (a) and (b) use random node selection, (c) and (d) rely on targeted selection. Networks in (a) and (c) contain $1,000$ nodes, while those in (b) and (d) consist of $10,000$ nodes.
Figure 20: Mean ECS using M-NMF method over the percentage of selected nodes on LFR benchmark graphs with $32$-dimensional embeddings. (a) and (b) implement random node selection, while (c) and (d) apply targeted selection; (a) and (c) correspond to networks with $1,000$ nodes, and (b) and (d) to networks with $10,000$ nodes.
Figure 21: Mean ECS using DeepWalk method over the percentage of selected nodes on LFR benchmark graphs with $32$-dimensional embeddings. Random selection is employed in (a) and (b), while targeted selection is used in (c) and (d). (a) and (c) represent networks with $1,000$ nodes, and (b) and (d) represent those with $10,000$ nodes.
Figure 22: Mean ECS using LINE method over the percentage of selected nodes on LFR benchmark graphs with $32$-dimensional embeddings. (a) and (b) follow random node selection, whereas (c) and (d) adopt targeted node selection. Networks with $1,000$ nodes correspond to (a) and (c), and those with $10,000$ nodes correspond to (b) and (d).
Figure 23: Mean ECS using node2vec method over the percentage of selected nodes on LFR benchmark graphs with $32$-dimensional embeddings. Random node selection is applied in (a) and (b), while (c) and (d) use targeted selection. (a) and (c) deal with $1,000$-node networks, and (b) and (d) with $10,000$-node networks.
Figure 24: Mean ECS for various methods over the percentage of selected nodes on the two real-world networks with $32$-dimensional embeddings. (a) and (b) use random node selection while (c) and (d) use targeted node selection. (a) and (c) corresponds to the email-EU-core network while (b) and (d) corresponds to the AS network.
Appendix B LINE method with higher embedding dimensions
For the LINE embedding method, our analysis of networks with 1,000 nodes in the first column of Fig. 12 reveals a distinct trend: the curves exhibit a steep decline followed by a subsequent increase. According to our experiments, increasing the embedding dimensions in LINE is expected to mitigate the sharpness of this decline and increase, rendering the pattern less pronounced, as indicated in Fig. 25. However, for other embedding methods, a higher embedding dimension doesn’t necessarily lead to better performance with higher similarity scores. For instance, M-NMF with higher dimensions results in lower ECS scores, as shown in Fig. 26, compared to the first column of Fig. 8.
Figure 25: Mean ECS using LINE method over the percentage of selected nodes uniformly at random on LFR benchmark graphs with $1,000$ nodes. The first plot displays results using $128$-dimensional embeddings; the second plot displays results using $256$-dimensional embeddings.
Figure 26: Mean ECS using M-NMF method over the percentage of selected nodes uniformly at random on LFR benchmark graphs with $1,000$ nodes. The first plot displays results using $128$-dimensional embeddings; the second plot displays results using $256$-dimensional embeddings.
@@ Acknowledgements.
The work of Zhi-Feng Wei is supported by the Department of Energy (DOE) Office of Advanced Scientific Computing Research (ASCR) through the ASCR Distinguished Computational Mathematics Postdoctoral Fellowship. Pacific Northwest National Laboratory (PNNL) is a multi-program national laboratory operated for the U.S. Department of Energy (DOE) by Battelle Memorial Institute under Contract No. DE-AC05-76RL01830. The research of Zhi-Feng Wei was supported in part by an appointment with the National Science Foundation (NSF) Mathematical Sciences Graduate Internship (MSGI) Program. This program is administered by the Oak Ridge Institute for Science and Education (ORISE) through an interagency agreement between the U.S. Department of Energy (DOE) and NSF. ORISE is managed for DOE by ORAU. All opinions expressed in this paper are the author’s and do not necessarily reflect the policies and views of NSF, ORAU/ORISE, or DOE. The research of Zhi-Feng Wei was supported in part by an appointment to the Oak Ridge National Laboratory GRO Program, sponsored by the U.S. Department of Energy and administered by the Oak Ridge Institute for Science and Education. This paper has been coauthored by UT-Battelle, LLC under Contract No. DE-AC05-00OR22725 with the U.S. Department of Energy. The publisher, by accepting the article for publication, acknowledges that the U.S. government retains a nonexclusive, paid up, irrevocable, world-wide license to publish or reproduce the published form of the manuscript, or allow others to do so, for U.S. government purposes. The DOE will provide public access to these results in accordance with the DOE Public Access Plan (http://energy.gov/downloads/doe-public-access-plan). This research was supported in part by Lilly Endowment, Inc., through its support for the Indiana University Pervasive Technology Institute.